UT 26 Aug 2020 - Night #5 of Restart

OSA: S. Allanson

ISp: E. Solheid

Mtn Manager: L. Durham

ISA: O. Kuhn (Tucson)

PROBLEMS

  1. Lost at least an hour or more due to /newdata running out of space. For future, even during restart, IA2 is strongly encouraged not to move data in/out of /newdata during nighttime operations.
  2. Spot seen on LBC-R (r-sloan) but seems to have vanished during skyflats - check filters or lenses?
  3. IT 6386 - MODS2B Sequencer - all sievesnaps at start of night affected, but seems to have abated durng on-sky - need to check exposures to verify it did not return. Has occurred of night each night.

PROPOSED PLAN OF ACTION:

  • MODS
    • Basic Checkout
      • Observe a standard star in all modes (30-45 min for 3 images).
        • Repeat if clear, otherwise take specstd for any modes used during the night
      • Dice-5 imaging offsets
        • Dumbell Nebula (first half)
        • M76 (2nd half)
    • Test Science OB
      • Sidereal
      • Non-Sidereal
    • Spectrophotometric standards (only if clear)
  • LBC- Red
    • DOFPIA / doHybrid - colliimatie on sky
    • Test Images
    • Rotator Center
  • Mixed mode MODS-1/LBC-Red tests
  • LUCI-2 (if released for on-sky)
    • Dectector Rotator
    • Rotator ZP (Longslit method - see OT and J.Power's notes)
    • Imaging Dither Test

Summary:

We obtained sufficient data to release MODS for science and got a start with LBCR.

We started with binocular MODS, and:
  • observed a spectrophotometric standard,
  • a science-like imaging sequence,
    • with async 45", 45" dithers (which required turning off the copointing check but ran well, save for one trailed pair of images
    • with sync offsets instead (which failed on at least two occasions when the asynchronicity between the 2 MODS exceeded the 60-sec timeout; see existing IT 6939. These observations only confirmed the recommendation not to use syncoffsets in imaging scripts.
  • ran a dithering sequence with the pointing mask to check the alignment of the DETXY coordinate system and mask
  • ran a long-slit observation (of PN NGC40).
  • observed a NS target
  • and then, as it appeared clear on the all-sky and the guide flux was stable (but note that there was significant haze! not to mention the many clouds earlier in the night), we
    • observed an photometric field in dual, blue and red-only imaging modes
    • observed a second standard star in all modes
Around 10:30 UT, we reconfigured MODS2-> LBCR and started to checkout LBCR with:
  • superfoc sequences
  • data to measure the rotator center
  • a science-like OB that dithered and guided

Issues:

We had a few issues:

MODS2B --- the first 7(!) sieveSnaps showed evidence of sequencer issues, but the remaining 80+ images were all good. IT 6836

newdata -- filled up during the night, a consequence of the work to resolve data flow to Repository (IT 8160)

The LBCR images through at least 3 filters showed a spot near the rotator center, but when I took sky flats through these same filters, and more, I saw no such spot. Was it a transient moth on the dewar window that flew away when we slewed? We observed 3 fields - I'll have to compare, but the sky flats show no evidence of dust or debris on the dewar or filters used (R, r, i, z, Y, F).

Log:

Times are in UT.

01:35 Ran up both MODS, woke them and took sieveSnap images. modsNc.20200826

First 5 mods2b images affected by sequencer issues. 0/5 good images.

01:45 Powered off/on HEB and did a quit,"ic" on m2.bc

01:50 Another two sieveSnaps 0006 and 0007 Again the mods2b images are bad.

02:09 Twilight sky flats on blank field 16+55:
  • Dual
  • 1K test, 0008. 1-sec at g and r. m1r: 15k, m1b
  • execMODS --modsN -e gr_skyflat_obs.txt
    • 3K @ gr 0009-13
  • execMODS --modsN -e gi(2)_skyflat_obs.txt
    • 3K @ gi 0014-
    • some stars are appearing on this set, but the flats are dithered.
  • Note - on the flats, mods2b is not showing the sequencer issue, but it can be hard to pull out of the flat field as we have seen recently for mods1r.
02:33 Going to the spectrophotometric standard, BD+33
  • Collimation/focus:
    • GCS is unable to write to Repository, IT 8169
    • Steve called Igor who is calling David E. They are working on it.
03:19 Steve is closing, since the 25-mic particle count went up above the closing limit.

03:35 Igor called back and said that David was able to reset permissions. We are testing it.
  • First, gcsclient readGuideCam left -e 1000
    • image appears in /lbt/data/share/tcs/lbto.20200826/left_guider00001.fits
    • but the image taken this way doesn't go to Repository
  • Second, Steve sent a test preset. It worked! David added a note to IT 8169 - sounds like he has resolved the problem for now, but more investigation is needed into why the NFS mounted Repository.
03:40 Steve is opening again. There are clouds, but all but sub-micron particulate counts are below their respective thresholds.

03:54 Back to BD+33:
  • Collimation/focus
  • acqBinoMODS bd332642.acq (19)
  • offsets made:
    • mods1: -0.699 11.748 rel
    • mods2: 3.434 8.282 rel
  • confirmatory thru-slit (20) look ok but a tad to the right
    • tweak using modsAlign - r on 0020
      • mods1: -0.312 0.104
      • mods2: -0223 -0.295
    • confirmatory thru-slit (21) looks better, but still the star is a bit right of center. Guide star is faint, ~15, and we have clouds and moonlight - perhaps this is a guiding issue?
  • execBInoMODS bd332642.obs
    • We are looking through clouds. Seeing is good, 0.77" on the guider.
    • So far, so good with the MODS2B. The only bad images were the first 7, all sieveSnaps.
    • < ~09:54-09:55MST lost guiding. On Dual Prism.
    • 10:01 - lost guiding again.
05:19 5-pt dither imaging script with offsets test, active mode:
  • execBinoMODS Dumbell_Nebula_Img_active.txt
    • 0.5" seeing
    • Error copointing limit 45", 45". Aborted.
    • Steve disabled the copointing check and we are repeating the script. It is working fine.
    • clouds causing the guide star to fade in and out. 05:45
      • mods1: after one offset, the offsetxy command timed out after 60-sec. Retry resumed but next image is streaked, mods1r 0039 and mods1b 0036.
      • mods2, at same time, slightly different error msg: no guide star found. Retry was ok for that side.
05:45 Repeat script but this time with syncoffset before each offset command. Let's see how things go as the 2 MODS get out of sync... We got a timeout error

execBinoMODS Dumbell_Nebula_Imct_sync.txt

MODS2 errors: bc syncoffset offsetxy 45 -45 absError: Command 'OFFSETXY 45 -45 abs' timed out after 60 secAbort, Retry, or Ignore? rRetrying command 'OFFSETXY 45 -45 absOFFSETXY 45 -45 abs

But I think that led to a stray sync flag

We had another timeout error on MODS2 on the offsetxy -45 45 abs

06:22 Longslit observation of PN NGC40
  • acquisition
    • MODS1: 49 & 50, offsetxy -0.606 10.970 rel
    • MODS2: 50 & 51, offsetxy 3.913 7.314 rel
  • execBinoMODS NGC40.obs
06:35 We ran out of disk space in newdata. Barry called Igor. Meanwhile, we continued taking spectra which appeared in modsNdata but only one (mods1r.20200826.0052.fits) in newdata.

06:55 Clouds - it's very overcast. Steve reports that he cannot see stars.

07:00 Igor made a new directory on Repository in which to temp store the files in newdata.

07:40 Newdata has more room now and we are moving on. NS target has a name that is not coming up in the NSIGUI (P29/Schwassmann-Wachmann 1), so we're doing something else while searching for alternate names.

07:45 M52 dither with pointing mask (to measure DETXY alignment with respect to mask alignment).

07:59 blue fitsflush

08:15 ~20 on each channel

08:15 NS test on Comet 29P/Schlassmann-Wachmann 1
  • execBInoMODS comet.img.txt
    • some scattered light from nearby star
    • guiding and WFSing is OK
08:45 Photometric Imaging field L92 in Strip82. It is not strictly photometric, but it is clear.
  • execBinoMODS L92dualim.txt
  • execBinoMODS L92redim.txt
  • execBinoMODS L92blueim.txt
09:20 Spectrophot Std Feige110
  • mods1r 86: -1.026 11.944
  • mods2r 90: 3.277 7.966
  • confirmatory images again to right. m1r by -0.16 and m2r by -0.4
  • observation: execBInoMODS feige110.obs
10:30-10:45 Reconfigured from MODS -> LBCR

10:45 Slewed to M76 for focus/collimation. That took a long time - 8 iterations. Seeing is very good and stable, but there was coma (z8) that WRS did not remove and the first 2 FPIA iterations were on focus only, giving -15276, then 9788. Then it added 2344 nm z8.

11:12 Superfoc
  • PA=0
  • PA=180
  • PA=0
  • Near the rotator center, there is a dust spot. (r-SLOAN)
11:29 Rotator Center measurement at 1h50d field
  • PA=0, 180, 90 and then 270
11:40 Science-like exposure
  • ptautoadjust (I forgot to collimate, but we had not moved far and collimation looks good)
  • run series of dithered, guided exposures at z, i and R.
  • In all of the LBCR images so far, taken through 4 different filters (r, z, i and R), there appears to be a spot near the rotator center.
12:06 Sky flats (although we'll need to clean the dewar window for real flats)
  • i-SLOAN: skyB~16 - bit dark still.
    • starting these at 12:17 or so and using 3-sec. The i-SLOAN flats don't show the dark ring... 121654.
  • r-SLOAN 12:19 skyB~14.5, 1-sec. The spot is gone here too...
  • R-Bessell - one flat below saturation but the spot appears to be gone
  • Y - 10k counts at 12:30, no spots
  • F972N20 - newdata filled up again. I continued the script, but will have to look for the flats later or on the PC.
12:38 Steve is closing up.
  • I turned off the LBCR (LBCB was OFF)
  • I put both MODS to sleep.

-- BarryRothberg - 25 Aug 2020
Topic revision: r9 - 26 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback